# Assessing the stability of ultracompact spinning boson stars with nonlinear evolutions

### v1

@@ III Theory
where $V(|\varphi|^{2})$ denotes the scalar potential. Here, we focus on the solitonic potential given by
$V(|\varphi|^{2})=\mu^{2}|\varphi|^{2}\left(1-\frac{2|\varphi|^{2}}{\sigma_{0}}\right)^{2},$ (2)
@@
$\nabla_{\alpha}\nabla^{\alpha}\varphi$ $=-\varphi\frac{\partial V(|\varphi|^{2})}{\partial|\varphi|^{2}},$ (4)
@@
$\mathrm{d}s^{2}=f\mathrm{d}t^{2}+$ $lf^{-1}\{g\left(\mathrm{d}r^{2}+r^{2}\mathrm{d}\theta^{2}\right)$
@@ V.1 Code infrastructure
ExoZvezda utilizes fourth-order finite-differencing of the CCZ4 system [54], where we fix the damping parameters to $\kappa_{1}=0.1/M$, $\kappa_{2}=0$ and $\kappa_{3}=1$ (note that our $\kappa_{1}$ is defined as in Eq. (27) of Ref. [55]). We complement the CCZ4 evolution scheme with the moving puncture gauge [56, 57], as given by Eqs. (28)-(31) of Ref. [50]. In all of our simulations we set up a computational domain of length $\mu L=1024$ with 7 additional AMR levels and utilize a bitant symmetry along the $z$-direction (i.e. the spin axis) [^2] Bitant symmetry implies that even (odd) perturbations exclude odd (even) $l$-modes during the evolution.. Our AMR is set up in such a way that the star is completely covered by the finest level. We consider resolutions in the ranges of $\mu\mathrm{d}x\approx$ 0.08-0.25 on the finest level, and find that our diagnostics converge at least at third order (see Appendix A.1 for more details).
The GHC code is also based on fourth-order accurate finite difference spatial discretization and Runge-Kutta time stepping, but utilizes the generalized harmonic formulation of the Einstein equations [53]. Throughout, we stick to stationary gauge, which fixes the source functions $H^{a}=\Box x^{a}$ to be constant in time and equal to the those of the stationary boson star solution in the form (III). Spatial infinity is included in the computational domain through a compactification of the Cartesian coordinates. Constraint damping terms, with $\kappa=2/M$ and $\rho=-0.5$, are used [58, 59]. For the purposes of this work, in 3+1 (2+1) evolutions we employ fixed mesh refinement with 7 (8) levels centered on the star. We note that in Table 1 the setup of the GHC code utilizing the generalized harmonic formulation results in gauge instabilities (we comment on this in more detail in Section VI.2.1 and Appendix C). The timescale of this gauge instability, however, is so long, that it is inaccessible by the 3+1 evolutions. We therefore additionally performed 2+1 evolutions with generalized harmonic formulation to investigate this further.
@@ V.2 Diagnostics
$\Phi_{\tilde{m}}=\int_{\mathcal{V}}\Phi e^{i\tilde{m}\phi}\mathrm{d}^{3}x,$ (10)
where $\Phi_{\tilde{m}}$ encodes the power of the $\tilde{m}-$th mode [^3] We use $\tilde{m}$ here to differentiate it from $m=1$ that enters the scalar field ansatz in Eq. (6). and we monitor the evolution of each $\Phi_{\tilde{m}}$ up to $\tilde{m}\leq 20$. Equation (10) is a common diagnostic for quantifying the nature of non-axisymmetric instabilities for neutron stars and BSs [60, 61, 62, 63]. The discretization error (especially for low resolution runs) results in a slight linear drift across the grid of the BS at later times. To account for this drift in the computation of $\Phi_{\tilde{m}}$ modes in Eq. (10), we recompute the azimuthal angle $\phi$ with the center-of-mass coordinates of the star, which we track by fitting a Gaussian to the conformal factor. Lastly, both codes utilize slightly different diagnostics to track the Einstein constraint violations. In particular, in ExoZvezda we track volume-weighted $L^{2}$-norms of the Hamiltonian and momentum constraint violations [see e.g. Eqs. (33-34) of Ref. [50] for their explicit definitions].
@@ VI.1 Main stability results
Here $\Delta$ denotes the width of the Gaussian profile and, for simplicity, we fix $\mu\Delta=1$. In this case the perturbation generally results in an excited star that starts to oscillate radially around its equilibrium value. These induced oscillations are long-lived, as can particularly be seen from $\Phi$ in the the case of larger amplitude $\epsilon=10^{-3}$ (see the right panel of Fig. 2). Analogous to the previous perturbation, BH formation occurs only for sufficiently large amplitude ($\epsilon\geq 0.01$ in this case), and the amplitude of the initial perturbation of S005, as measured by $\Phi$, is orders of magnitude larger between the cases with explicit perturbation (14) and the one without, as can be seen by comparing Figs. 2 and 1.
@@ VI.2.2 Constraint damping
As first pointed out in Ref. [16], we have also found that constraint damping plays a key role in long-term evolutions using the ExoZvezda code. In particular, in addition to the results presented above obtained using the CCZ4 formulation, we evolved the S005 model in the exact same set-up of Table 1 (and identical gauge), but with no constraint damping using BSSNOK formulation [66, 67, 68]. All of our BSSNOK simulations, however, suffered from a gradual growth in constraint violations, as illustrated in Fig. 4, reminiscent of constraint growth reported in neutron stars evolutions in Ref. [69] (see their Fig. 5). The BSSNOK constraint subsystem has previously been shown to possess a zero-speed characteristic (see e.g. Section VI of Ref. [70] or Section II of Ref. [71]), which can accumulate in time and lead to large constraint violations during the evolution. Even at the initial time we inevitably accrue some constraint violations due to interpolation errors and imperfections in initial data [^12] We note that we achieve first order point-wise convergence of the Hamiltonian constraint at the initial time slice.; particularly around the regions of the star where the scalar field gradients are large. In BSSNOK simulations, however, these constraint violations remain stationary on the grid and gradually start to accumulate in time, whilst in CCZ4 such constraint violating modes propagate away quickly. In fact, in BSSNOK runs we observe a deterioration in the convergence of the $L^{2}$-norm of the constraints to only first order during the evolution. Furthermore, CCZ4 reduces the Hamiltonian constraint violations by two orders of magnitude and the drift in $\Phi_{\rm{max}}$ by one order of magnitude.
@@ Acknowledgements.
T.E. thanks Ulrich Sperhake, Seppe Staelens, and Gareth Marks for useful discussions. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. We acknowledge support by the Cambridge Service for Data Driven Discovery at the University of Cambridge and Durham University through DiRAC Projects ACTP284 and ACTP and STFC capital Grants No. ST/P002307/1 and No. ST/R002452/1, and STFC operations Grant No. ST/R00689X/1. W.E. acknowledges support from a Natural Sciences and Engineering Research Council of Canada Discovery Grant and an Ontario Ministry of Colleges and Universities Early Researcher Award. This research was enabled in part by support provided by SciNet (www.scinethpc.ca) and the Digital Research Alliance of Canada (al- liancecan.ca). Computations were done on the CSD3 (Cambridge), Cosma7 and 8 (Durham) and Niagara (University of Toronto) clusters.
@@ Appendix D List of evolved models
In Table 2, we summarize all the models (including the S005 and S02 featured in the main text) that we have investigated for stability.

### v2

@@ III Theory
where $V(|\varphi|^{2})$ denotes the scalar potential. Here, we focus on the solitonic potential given by [^1] We note that our convention for the action is different from that of Ref. [60]. In particular, what we call $\sigma_{0}$ here would be called $\sigma_{0}/\sqrt{2}$ in the conventions of Ref. [60].
$V(|\varphi|^{2})=\mu^{2}|\varphi|^{2}\left(1-\frac{2|\varphi|^{2}}{\sigma_{0}^{2}}\right)^{2},$ (2)
@@
$\nabla_{\alpha}\nabla^{\alpha}\varphi$ $=\varphi\frac{\partial V(|\varphi|^{2})}{\partial|\varphi|^{2}},$ (4)
@@
$\mathrm{d}s^{2}=-f\mathrm{d}t^{2}+$ $lf^{-1}\{g\left(\mathrm{d}r^{2}+r^{2}\mathrm{d}\theta^{2}\right)$
@@ V.1 Code infrastructure
ExoZvezda utilizes fourth-order finite-differencing of the CCZ4 system [9], where we fix the damping parameters to $\kappa_{1}=0.1/M$, $\kappa_{2}=0$ and $\kappa_{3}=1$ (note that our $\kappa_{1}$ is defined as in Eq. (27) of Ref. [10]). We complement the CCZ4 evolution scheme with the moving puncture gauge [25, 12], as given by Eqs. (28)-(31) of Ref. [66]. We further utilize sixth-order Kreiss-Oliger dissipation with a dissipation parameter set to 0.3 (see e.g. Section 2.2.4. of Ref. [30]), although find our conclusions unchanged for lower values. In all of our simulations we set up a computational domain of length $\mu L=1024$ with 7 additional AMR levels and utilize a bitant symmetry along the $z$-direction (i.e. the spin axis) [^3] Bitant symmetry implies that even (odd) perturbations exclude odd (even) $l$-modes during the evolution.. Our AMR is set up in such a way that the star is completely covered by the finest level. We consider resolutions in the ranges of $\mu\mathrm{d}x\approx$ 0.08-0.25 on the finest level, and find that our diagnostics converge at least at third order (see Appendix A.1 for more details).
The GHC code is also based on fourth-order accurate finite difference spatial discretization and Runge-Kutta time stepping, with sixth-order Kreiss-Oliger dissipation, but utilizes the generalized harmonic formulation of the Einstein equations [64]. Throughout, we stick to stationary gauge, which fixes the source functions $H^{a}=\Box x^{a}$ to be constant in time and equal to the those of the stationary boson star solution in the form (III). Spatial infinity is included in the computational domain through a compactification of the Cartesian coordinates. Constraint damping terms, with $\kappa=2/M$ and $\rho=-0.5$, are used [46, 65]. For the purposes of this work, in 3+1 (2+1) evolutions we employ fixed mesh refinement with 7 (8) levels centered on the star. We note that in Table 1 the setup of the GHC code utilizing the generalized harmonic formulation results in gauge instabilities (we comment on this in more detail in Section VI.2.1 and Appendix C). The timescale of this gauge instability, however, is so long, that it is inaccessible by the 3+1 evolutions. We therefore additionally performed 2+1 evolutions with generalized harmonic formulation to investigate this further.
@@ V.2 Diagnostics
$\Phi_{\tilde{m}}=\int_{\mathcal{V}}|\partial_{t}|\varphi|^{2}|e^{i\tilde{m}\phi}\mathrm{d}^{3}x,$ (10)
where $\Phi_{\tilde{m}}$ encodes the power of the $\tilde{m}-$th mode [^4] We use $\tilde{m}$ here to differentiate it from $m$ that enters the scalar field ansatz in Eq. (6). and we monitor the evolution of each $\Phi_{\tilde{m}}$ up to $\tilde{m}\leq 20$. Equation (10) is a common diagnostic for quantifying the nature of non-axisymmetric instabilities for neutron stars and BSs [51, 37, 39, 74]. The discretization error (especially for low resolution runs) results in a slight linear drift across the grid of the BS at later times. To account for this drift in the computation of $\Phi_{\tilde{m}}$ modes in Eq. (10), we recompute the azimuthal angle $\phi$ with the center-of-mass coordinates of the star, which we track by fitting a Gaussian to the conformal factor. Lastly, both codes utilize slightly different diagnostics to track the Einstein constraint violations. In particular, in ExoZvezda we track volume-weighted $L^{2}$-norms of the Hamiltonian and momentum constraint violations [see e.g. Eqs. (33-34) of Ref. [66] for their explicit definitions].
@@ VI.1 Main stability results
Here $\Delta$ denotes the width of the Gaussian profile and, for simplicity, we fix $\mu\Delta=1$. In this case the perturbation generally results in an excited star that starts to oscillate radially around its equilibrium state. These induced oscillations are long-lived, as can particularly be seen from $\Phi$ in the the case of larger amplitude $\epsilon=10^{-3}$ (see the right panel of Fig. 2). Analogous to the previous perturbation, BH formation occurs only for sufficiently large amplitude ($\epsilon\geq 0.01$ in this case), and the amplitude of the initial perturbation of S005, as measured by $\Phi$, is orders of magnitude larger between the cases with explicit perturbation (14) and the one without, as can be seen by comparing Figs. 2 and 1.
@@ VI.2.2 Constraint damping
As first pointed out in Ref. [60], we have also found that constraint damping plays a key role in long-term evolutions using the ExoZvezda code. In particular, in addition to the results presented above obtained using the CCZ4 formulation, we evolved the S005 model in the exact same set-up of Table 1 (and identical gauge), but with no constraint damping using the BSSNOK formulation [16, 71, 63]. All of our BSSNOK simulations, however, suffered from a gradual growth in constraint violations, as illustrated in Fig. 4, reminiscent of constraint growth reported in neutron stars evolutions in Ref. [19] (see their Fig. 5). The BSSNOK constraint subsystem has previously been shown to possess a zero-speed characteristic (see e.g. Section VI of Ref. [20] or Section II of Ref. [23]), which can accumulate in time and lead to large constraint violations during the evolution. Even at the initial time we inevitably accrue some constraint violations due to interpolation errors and imperfections in initial data [^13] We note that we achieve first order point-wise convergence of the Hamiltonian constraint at the initial time slice.; particularly around the regions of the star where the scalar field gradients are large. In BSSNOK simulations, however, these constraint violations remain stationary on the grid and gradually start to accumulate in time, whilst in CCZ4 such constraint violating modes propagate away quickly. In fact, in BSSNOK runs we observe a deterioration in the convergence of the $L^{2}$-norm of the constraints to only first order during the evolution. Furthermore, CCZ4 reduces the Hamiltonian constraint violations by two orders of magnitude and the drift in $\Phi_{\rm{max}}$ by one order of magnitude.
@@ Acknowledgements.
T.E. thanks Ulrich Sperhake, Seppe Staelens, and Gareth Marks for useful discussions. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. We acknowledge support by the Cambridge Service for Data Driven Discovery at the University of Cambridge and Durham University through DiRAC Projects ACTP284 and ACTP and STFC capital Grants No. ST/P002307/1 and No. ST/R002452/1, and STFC operations Grant No. ST/R00689X/1. W.E. acknowledges support from a Natural Sciences and Engineering Research Council of Canada Discovery Grant and an Ontario Ministry of Colleges and Universities Early Researcher Award. This research was enabled in part by support provided by SciNet (www.scinethpc.ca) and the Digital Research Alliance of Canada (alliancecan.ca). Computations were done on the CSD3 (Cambridge), Cosma7 and 8 (Durham) and Niagara (University of Toronto) clusters.
@@ Appendix D List of evolved models
In Table 2, we summarize all the models (including the S005 and S02 featured in the main text) that we have investigated for stability. Further, in Fig. 9 we display where in the parameter space these models reside.
@@
Figure 9: Main properties (ADM mass and BS frequency) of $\sigma_{0}=0.05,\,m=1$; $\sigma_{0}=0.05,\,m=0$; and $\sigma_{0}=0.2,\,m=3$ BS families featured in the Table 2. We use the star markers to indicate the S005 and S02 solutions discussed in the main text.